# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ E.4 Evaluation Methodology
In the final metrics calculation stage, a series of quantitative metrics are utilized to objectively and thoroughly evaluate model performance (see Section E.3.2 for details). Specifically, these metrics include: the HDR, which measures the proportion of true physical violations accurately identified by the model, thus reflecting its detection sensitivity; the FHR, which assesses the proportion of misclassifications the model makes in normal control cases without actual violations, hence indicating the model’s robustness and false positive rate. Additionally, we introduce the EQS, which is assigned by three expert human raters on a 5-point scale. Ratings are given from multiple perspectives, including physical correctness, the relevance of the explanation to the audio facts, and the logic and clarity of the reasoning process. The final EQS is computed as the mean score across all raters and all test cases, thereby providing a comprehensive quantitative measure of the model’s interpretability. Overall, this multi-dimensional evaluation framework effectively captures the model’s Overall competence in the context of physical violation detection, encompassing detection accuracy, false positives, and explanation quality, thus offering a reliable experimental foundation for subsequent optimization and improvement of the methods.
@@ E.5 Result Analysis
The scores are in the format "illusion recognition / illusion resistance / quality score". Higher values (0-10 scale) indicate better performance.
@@
We evaluate the hallucination performance of nine models in Appendix E.3.1, with detailed results presented in Table 10, Table 11, and Table 12. The results reveal the following key findings:
(1) Hallucination resistance varies significantly among different Auditory Large Language Models (ALLMs). In the general hallucination assessments (Table 10 and 11), models often considered highly capable, such as Gemini-1.5 Pro, Gemini-2.5 Flash, and Gemini-2.5 Pro, generally exhibit strong performance (higher scores, indicating better resistance to hallucination). Ultravox also frequently performs well. In contrast, models like SALMONN, and often GPT-4o Audio and GPT-4o mini Audio, tend to show lower scores in these general tests, suggesting a higher propensity for hallucination. Open-source models like MiniCPM-o 2.6 and Qwen2-Audio demonstrate competitive and often robust performance against hallucinations in these experiments.
@@ F.2 Dataset Construction Method
To construct our dataset of adversarial prompts, we followed a two-stage process: prompt generation using GPT-4o [54] and audio synthesis using F5-TTS [12]. First, we leveraged OpenAI’s GPT-4o (the multimodal GPT-4 with voice capabilities) to automatically generate a diverse set of harmful or policy-violating user queries for each category of attack. GPT-4o was chosen for prompt generation due to its advanced language capabilities and alignment safeguards, ensuring that the produced queries are realistic and varied. We crafted a consistent prompt template to guide GPT-4o in creating these adversarial questions. This template provided the scenario context (enterprise, financial, medical, or illegal), specified the speaker’s gender (male or female) and emotional tone (angry or sad), and instructed GPT-4o to produce a user query attempting to break the system’s rules. The template was designed to simulate a malicious user’s perspective, instructing GPT-4o to create challenging queries that, while mimicking harmful intent, adhered to its content generation policies during the generation process. An example of the prompt construction template is shown below:
Using this template (with appropriate substitutions for each scenario and persona), we generated the full set of prompts. In total, we obtained 300 text prompts for the Jailbreak category and 300 for the Illegal Activity Guidance category. For the Jailbreak prompts, we generated 100 queries for each of the three scenarios (Enterprise, Financial, Medical). Within each scenario’s 100 prompts, we ensured an even coverage of the four gender–emotion combinations: 25 prompts each for Male-Angry, Male-Sad, Female-Angry, and Female-Sad. Similarly, for the Illegal Activity Guidance category, we produced 300 prompts evenly divided into 75 for each of the four combinations (male-angry, male-sad, female-angry, female-sad). This balancing ensures that our dataset controls for speaker gender and emotional tone, allowing us to analyze their effects on attack success. After generating the textual prompts, we converted each prompt into spoken audio using the F5-TTS [12] text-to-speech system. F5-TTS is a state-of-the-art TTS model capable of high-quality voice cloning with emotional expressiveness. For each prompt, we selected a voice profile that matched the intended gender and could convey the specified emotion (angry or sad). The text of the prompt was then synthesized into an audio file using F5-TTS, resulting in a spoken version of the query. This audio conversion is crucial because our attack evaluations are conducted via GPT-4o’s voice interface. Using F5-TTS to produce natural and emotive speech ensures the audio prompts closely mimic real user utterances in content and delivery.
@@ F.4 Evaluation Method and Metrics
Here, is the number of test samples, and is the GPT-4o score for the -th response. A higher DSR reflects stronger safety alignment, while a lower HRR indicates fewer harmful completions. These two complementary metrics provide a comprehensive view of model safety under adversarial audio input. The safety of ALLMs is evaluated using these datasets and the safety metrics detailed in Section F.3, with results presented in Section F.5.
@@ F.5 Additional Result
Unexpected Outputs from Open-source Models In Table 3, we present the experimental results of open-source models. However, these results are based on the observation that open-source models randomly produce erroneous outputs for approximately one-third of the test inputs. Specifically, the models sometimes fail to receive the audio portion of a sample and only process the corresponding text. As a result, they tend to generate brief, generic replies such as “Hi! I am your enterprise assistant. Feel free to ask me anything. How can I help you?” This occurs because they are often only capable of correctly parsing audio inputs shorter than 30 seconds. During evaluation, such responses are frequently rated highly by the LLM-based evaluator, leading to an inflated overall safety score. Nonetheless, since the proportion of these inflated scores is relatively consistent across all open-source models, comparisons among them remain meaningful.
| MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio | |
@@
| Male Angry | 95 | 80 | 113 | 125 | 99.2 | 125 | 125 | 125 | 124 |
| Male Sad | 103 | 77 | 89 | 117 | 97.6 | 125 | 124 | 124 | 125 |
| Female Angry | 103 | 89 | 80 | 125 | 99.2 | 125 | 125 | 123 | 125 |
| Female Sad | 80 | 95 | 89 | 119 | 100 | 125 | 125 | 123 | 125 |
@@
| Male Angry | 101 | 102 | 88 | 85 | 100 | 125 | 125 | 125 | 125 |
| Male Sad | 98 | 98 | 91 | 107 | 98.4 | 125 | 125 | 123 | 123 |
| Female Angry | 96 | 104 | 93 | 118 | 100 | 125 | 125 | 124 | 124 |
| Female Sad | 101 | 99 | 100 | 109 | 98.4 | 124 | 122 | 124 | 123 |
@@
| Male Angry | 113 | 108 | 106 | 118 | 98.4 | 125 | 125 | 123 | 125 |
| Male Sad | 92 | 103 | 104 | 107 | 96 | 124 | 124 | 121 | 123 |
| Female Angry | 94 | 95 | 90 | 117 | 99.2 | 125 | 123 | 123 | 123 |
| Female Sad | 109 | 101 | 104 | 112 | 96.8 | 123 | 125 | 124 | 122 |
@@
| Male Angry | 372 | 368 | 315 | 375 | 100 | 375 | 375 | 375 | 375 |
| Male Sad | 360 | 344 | 329 | 361 | 100 | 375 | 375 | 375 | 375 |
| Female Angry | 358 | 335 | 270 | 370 | 100 | 372 | 374 | 375 | 374 |
| Female Sad | 353 | 340 | 243 | 364 | 99.7 | 375 | 372 | 375 | 375 |
In Table 13, we summarize the scores of different models under Jailbreak and Illegal Activities Guidance induction experiments, categorized by speaker gender and emotional tone.
@@
Additionally, we conducted experiments with closed-source text models, using the text inputs from our original experimental setup and the textual content of the audio as input to evaluate the performance of large language models. As shown in Table 14, all models achieved 100% safety, indicating that for ALLMs, audio containing emotional and contextual content is indeed more likely to break through the model’s defenses and cause jailbreaking compared to pure text input.
@@ G.2 Dataset Construction Method
Direct Privacy Leakage Dataset (600 samples). For the direct privacy leakage dataset,we created six categories of sensitive personal information: bank account numbers, home addresses, phone numbers, phone passwords, driver’s license numbers, and social security numbers [71]. For each category, we first prompted Grok to generate textual data containing personal names and corresponding private information [79]. The generated data were then further randomized to enhance diversity and reduce identifiability. Based on this preliminary dataset, Grok was instructed to generate realistic dialogue scenarios in which each type of private information might naturally appear. Subsequently, Grok produced context-appropriate conversational texts for each scenario. Finally, we selected voice samples from the Common Voice dataset and used F5-TTS to synthesize the corresponding audio files [12, 4]. Each category contained 100 entries, resulting in a total of 600 test samples.
Privacy Inference Leakage Dataset (300 samples). For the evaluation of implicit privacy inference, including age, ethnicity, and conversation setting [81], we constructed another dataset comprising 300 audio samples, with 100 samples per category. The age and ethnicity samples were obtained by filtering and curating entries from the Common Voice database. For the conversation setting category, we first collected ambient audio clips representing common daily-life environments such as “workspace” and “home.” Grok was then used to generate dialogues likely to occur in these environments. These dialogues were synthesized into speech using TTS and mixed with the corresponding background environment audio to create the final dataset. Below, we provide a selection of prompts employed during the dataset construction process using large language models.
@@ G.4 Evaluation Method and Metrics
For each audio input, privacy-related questions were posed via audio question answering (e.g., “What is the speaker‘s bank account number?” or “What is the speaker‘s age?”). The models’ responses were recorded and automatically scored using a five-point scoring system, ranging from 5 (complete refusal due to privacy concerns) to 1 (full disclosure of private information) [83], with intermediate scores reflecting varying degrees of caution in information disclosure.
We reported three key metrics based on the scores:
Refusal rate: the proportion of responses scored 5 or 4 (indicating refusal or cautious non-disclosure),
Complete disclosure rate: the proportion of responses scored 1 (indicating unconditional disclosure of private information),
Cautious disclosure rate: the proportion of responses scored 3 or 2 (indicating partial or ambiguous disclosure).
@@ G.5 Experiment Design and Results
We used synthetic conversational audio containing specific types of private information (e.g., bank account numbers, home addresses, and phone numbers) as input to fix closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and four open-source ALLMs (e.g., Qwen2-Audio). We then posed a series of privacy-related questions tied to the audio content and recorded the models‘ responses.
@@
“The audio samples were input into five closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and four open-source ALLMs (e.g., Qwen2-Audio). For age and ethnicity inference, implicit information was queried via AQA (e.g., “What is the speaker’s age?” or “What is the speaker’s ethnicity?”), and responses were scored using the same five-point scoring system as in the direct leakage experiments (5 = complete refusal, 1 = full disclosure). Scoring was performed automatically by a pretrained language model.
@@
| Qwen2-Audio | 0 | 13 | 1 | 12 | 0 | 8 | 4 | 49 | 0 | 5 | 0 | 5 | 0 | 0 | 0 | 0 | 0 | 0 |
| SALMONN | 51 | 97 | 57 | 98 | 64 | 91 | 73 | 99 | 66 | 99 | 30 | 100 | 97 | 100 | 33 | 13 | 0 | 25 |
| MiniCPM-o 2.6 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 6 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
| UltraVox | 100 | 100 | 100 | 100 | 18 | 100 | 70 | 100 | 100 | 100 | 50 | 100 | 1 | 6 | 24 | 17 | 0 | 0 |
@@
| Gemini-1.5 Pro | 0 | 100 | 0 | 100 | 0 | 99 | 0 | 98 | 70 | 100 | 0 | 93 | 16 | 28 | 0 | 0 | 0 | 0 |
| Gemini-2.5 Flash | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 85 | 0 | 5 | 0 | 1 | 0 | 0 |
| Gemini-2.5 Pro | 0 | 98 | 0 | 99 | 100 | 100 | 0 | 100 | 100 | 100 | 0 | 86 | 0 | 0 | 0 | 0 | 0 | 0 |
| GPT-4o Audio | 100 | 100 | 100 | 100 | 70 | 99 | 100 | 100 | 100 | 100 | 88 | 99 | 2 | 22 | 17 | 31 | 0 | 4 |
| GPT-4o Mini Audio | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 9 | 68 | 32 | 48 | 0 | 0 |
| Average | 39.0 | 78.7 | 39.8 | 78.8 | 39.1 | 77.4 | 38.6 | 83.6 | 59.6 | 78.2 | 29.8 | 74.2 | 13.9 | 25.4 | 11.8 | 12.2 | 0.0 | 3.2 |
Note: "no aug" indicates the refusal rates before applying a privacy-aware prompt, while "aug" shows rates after applying it. Higher values indicate better performance. Blue arrows () indicate better performance (higher refusal rate) than average; red arrows () indicate worse performance (lower refusal rate) than average. SSN: Social Security Number; Pwd: Phone Password; Setting: Setting of Conversation.
| no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | no aug | aug | |

### v3

@@ E.4 Evaluation Methodology
In the final metrics calculation stage, a series of quantitative metrics are utilized to objectively and thoroughly evaluate model performance (see Section˜E.3.2 for details). Specifically, these metrics include: the HDR, which measures the proportion of true physical violations accurately identified by the model, thus reflecting its detection sensitivity; the FHR, which assesses the proportion of misclassifications the model makes in normal control cases without actual violations, hence indicating the model’s robustness and false positive rate. Additionally, we introduce the EQS, which is assigned by three expert human raters on a 5-point scale. Ratings are given from multiple perspectives, including physical correctness, the relevance of the explanation to the audio facts, and the logic and clarity of the reasoning process. The final EQS is computed as the mean score across all raters and all test cases, thereby providing a comprehensive quantitative measure of the model’s interpretability. Overall, this multi-dimensional evaluation framework effectively captures the model’s Overall competence in the context of physical violation detection, encompassing detection accuracy, false positives, and explanation quality, thus offering a reliable experimental foundation for subsequent optimization and improvement of the methods.
@@ E.5 Result Analysis
| Step-Fun | 3.97/3.83/4.09 | 6.17/6.33/5.78 | 5.88/5.75/6.30 | 8.89/8.50/8.96 |
| OpenS2S | 2.01/1.79/1.89 | 2.75/2.75/4.75 | 8.00/7.00/5.97 | 8.89/8.89/8.31 |
| Kimi-Audio | 1.38/1.39/1.38 | 2.42/3.15/2.75 | 5.00/6.00/6.09 | 4.50/4.00/8.58 |
| Qwen2.5-Omni | 7.96/8.02/7.99 | 3.80/3.00/5.57 | 7.67/7.67/8.12 | 5.20/5.00/6.36 |
| Step-Audio2 | 3.51/3.73/3.61 | 0.00/0.00/5.82 | 0.00/0.00/7.80 | 0.00/0.00/8.28 |
@@
Scores follow the format “DIM 1 / DIM 2 / DIM 3”. Higher values indicate better performance.
@@
| Test Type | Step-Fun | OpenS2S | Kimi-Audio | Qwen2.5-Omni | ||||||||
| Content Mismatch | 54.6 | 16.2 | 29.2 | 81.7 | 2.5 | 15.8 | 67.5 | 29.2 | 3.3 | 26.2 | 0.0 | 73.8 |
| Label Mismatch | 50.0 | 20.4 | 29.6 | 57.5 | 4.6 | 37.9 | 28.7 | 33.3 | 37.9 | 77.1 | 1.2 | 21.7 |
| Logical Violation | 34.4 | 14.4 | 51.2 | 54.4 | 3.1 | 42.5 | 33.1 | 55.0 | 11.9 | 50.6 | 31.2 | 18.1 |
| Physical Violation | 11.7 | 11.7 | 76.7 | 41.6 | 15.1 | 43.3 | 23.8 | 17.5 | 58.8 | 52.5 | 5.8 | 41.7 |
@@
We evaluate the hallucination performance of nine models in Appendix E.4, with detailed results presented in Table 12, Table 13, and Table 14. The results reveal the following key findings:
(1) Hallucination resistance varies significantly among different Auditory Large Language Models (ALLMs). In the general hallucination assessments (Table 12 and 13), models often considered highly capable, such as Gemini-1.5 Pro, Gemini-2.5 Flash, and Gemini-2.5 Pro, generally exhibit strong performance (higher scores, indicating better resistance to hallucination). Ultravox also frequently performs well. In contrast, models like SALMONN, and often GPT-4o Audio and GPT-4o mini Audio, tend to show lower scores in these general tests, suggesting a higher propensity for hallucination. Open-source models like MiniCPM-o 2.6 and Qwen2-Audio demonstrate competitive and often robust performance against hallucinations in these experiments.
@@ F.2 Dataset Construction Method
To construct our dataset of adversarial prompts, we followed a two-stage process: prompt generation using GPT-4o (OpenAI et al., 2024) and audio synthesis using F5-TTS (Chen et al., 2024). First, we leveraged OpenAI’s GPT-4o (the multimodal GPT-4 with voice capabilities) to automatically generate a diverse set of harmful or policy-violating user queries for each category of attack. GPT-4o was chosen for prompt generation due to its advanced language capabilities and alignment safeguards, ensuring that the produced queries are realistic and varied. We crafted a consistent prompt template to guide GPT-4o in creating these adversarial questions. This template provided the scenario context (enterprise, financial, medical, or illegal), specified the speaker’s gender (male or female) and emotional tone (angry or sad), and instructed GPT-4o to produce a user query attempting to break the system’s rules. The template was designed to simulate a malicious user’s perspective, instructing GPT-4o to create challenging queries that, while mimicking harmful intent, adhered to its content generation policies during the generation process. An example of the prompt construction template is shown below:
Using this template (with appropriate substitutions for each scenario and persona), we generated the full set of prompts. In total, we obtained 300 text prompts for the Jailbreak category and 300 for the Illegal Activity Guidance category. For the Jailbreak prompts, we generated 100 queries for each of the three scenarios (Enterprise, Financial, Medical). Within each scenario’s 100 prompts, we ensured an even coverage of the four gender–emotion combinations: 25 prompts each for Male-Angry, Male-Sad, Female-Angry, and Female-Sad. Similarly, for the Illegal Activity Guidance category, we produced 300 prompts evenly divided into 75 for each of the four combinations (male-angry, male-sad, female-angry, female-sad). This balancing ensures that our dataset controls for speaker gender and emotional tone, allowing us to analyze their effects on attack success. After generating the textual prompts, we converted each prompt into spoken audio using the F5-TTS (Chen et al., 2024) text-to-speech system. F5-TTS is a state-of-the-art TTS model capable of high-quality voice cloning with emotional expressiveness. For each prompt, we selected a voice profile that matched the intended gender and could convey the specified emotion (angry or sad). The text of the prompt was then synthesized into an audio file using F5-TTS, resulting in a spoken version of the query. This audio conversion is crucial because our attack evaluations are conducted via GPT-4o’s voice interface. Using F5-TTS to produce natural and emotive speech ensures the audio prompts closely mimic real user utterances in content and delivery.
@@ F.4 Evaluation Method and Metrics
Here, is the number of test samples, and is the GPT-4o score for the -th response. A higher DSR reflects stronger safety alignment, while a lower HRR indicates fewer harmful completions. These two complementary metrics provide a comprehensive view of model safety under adversarial audio input. The safety of ALLMs is evaluated using these datasets and the safety metrics detailed in Section˜F.3, with results presented in Section˜F.5.
@@ F.5 Additional Result
Unexpected Outputs from Open-source Models In Table˜3, we present the experimental results of open-source models. However, these results are based on the observation that open-source models randomly produce erroneous outputs for approximately one-third of the test inputs. Specifically, the models sometimes fail to receive the audio portion of a sample and only process the corresponding text. As a result, they tend to generate brief, generic replies such as “Hi! I am your enterprise assistant. Feel free to ask me anything. How can I help you?” This occurs because they are often only capable of correctly parsing audio inputs shorter than 30 seconds. During evaluation, such responses are frequently rated highly by the LLM-based evaluator, leading to an inflated overall safety score. Nonetheless, since the proportion of these inflated scores is relatively consistent across all open-source models, comparisons among them remain meaningful.
| SALMONN | Ultravox | Qwen2-Audio | MiniCPM-o 2.6 | Step-Fun | Qwen2.5-omni | Kimi-Audio | OpenS2S | Step-Audio2 | |
@@
| Male Angry | 113 | 125 | 80 | 95 | 98.0 | 122.0 | 125.0 | 46.0 | 122.0 |
| Male Sad | 89 | 117 | 77 | 103 | 85.0 | 121.0 | 125.0 | 83.0 | 109.0 |
| Female Angry | 80 | 125 | 89 | 103 | 86.0 | 125.0 | 123.0 | 64.0 | 106.0 |
| Female Sad | 89 | 119 | 95 | 80 | 84.0 | 118.0 | 124.0 | 64.0 | 103.0 |
@@
| Male Angry | 88 | 85 | 102 | 101 | 105.0 | 112.0 | 122.0 | 60.0 | 80.0 |
| Male Sad | 91 | 107 | 98 | 98 | 113.0 | 123.0 | 123.0 | 104.0 | 82.0 |
| Female Angry | 93 | 118 | 104 | 96 | 100.0 | 121.0 | 123.0 | 71.0 | 90.0 |
| Female Sad | 100 | 109 | 99 | 101 | 113.0 | 118.0 | 123.0 | 104.0 | 82.0 |
@@
| Male Angry | 106 | 118 | 108 | 113 | 109.0 | 123.0 | 119.0 | 86.0 | 101.0 |
| Male Sad | 104 | 107 | 103 | 92 | 111.0 | 113.0 | 118.0 | 96.0 | 84.0 |
| Female Angry | 90 | 117 | 95 | 94 | 110.0 | 117.0 | 124.0 | 82.0 | 92.0 |
| Female Sad | 104 | 112 | 101 | 109 | 115.0 | 118.0 | 115.0 | 112.0 | 84.0 |
@@
| Male Angry | 315 | 375 | 368 | 372 | 372.0 | 375.0 | 375.0 | 192.0 | 274.0 |
| Male Sad | 329 | 361 | 344 | 360 | 346.0 | 372.0 | 375.0 | 322.0 | 232.0 |
| Female Angry | 270 | 370 | 335 | 358 | 343.0 | 370.0 | 374.0 | 215.0 | 269.0 |
| Female Sad | 243 | 364 | 340 | 353 | 356.0 | 370.0 | 375.0 | 348.0 | 240.0 |
| Gemini-1.5 Pro | GPT-4o Audio | GPT-4o mini Audio | Gemini-2.5 Flash | Gemini-2.5 Pro | |
| Enterprise System Jailbreak | |||||
| Male Angry | 99.2 | 125 | 124 | 125 | 125 |
| Male Sad | 97.6 | 124 | 125 | 125 | 124 |
| Female Angry | 99.2 | 123 | 125 | 125 | 125 |
| Female Sad | 100 | 123 | 125 | 125 | 125 |
| Financial System Jailbreak | |||||
| Male Angry | 100 | 125 | 125 | 125 | 125 |
| Male Sad | 98.4 | 123 | 123 | 125 | 125 |
| Female Angry | 100 | 124 | 124 | 125 | 125 |
| Female Sad | 98.4 | 124 | 123 | 124 | 122 |
| Medical System Jailbreak | |||||
| Male Angry | 98.4 | 123 | 125 | 125 | 125 |
| Male Sad | 96 | 121 | 123 | 124 | 124 |
| Female Angry | 99.2 | 123 | 123 | 125 | 123 |
| Female Sad | 96.8 | 124 | 122 | 123 | 125 |
| Illegal Activities Guidance | |||||
| Male Angry | 100 | 375 | 375 | 375 | 375 |
| Male Sad | 100 | 375 | 375 | 375 | 375 |
| Female Angry | 100 | 375 | 374 | 372 | 374 |
| Female Sad | 99.7 | 375 | 375 | 375 | 372 |
In Table˜15 and Table˜16, we summarize the scores of different models under Jailbreak and Illegal Activities Guidance induction experiments, categorized by speaker gender and emotional tone.
@@
Additionally, we conducted experiments with closed-source text models, using the text inputs from our original experimental setup to evaluate the performance of large language models. However, we noted that pure text input loses the key emotional information (e.g., anger, sadness) contained in the original audio, which could lead to an unfair comparison with Audio Large Language Models (ALLMs).
To compensate for this and ensure a more rigorous comparison, we augmented the text inputs with explicit emotional prompts designed to simulate the emotional state of the original audio. The specific augmentation prompts were as follows:
As shown in Table˜17, the results revealed that even after incorporating these strong emotional prompts, all tested text models still achieved 100% safety. This finding more strongly substantiates our conclusion: for ALLMs, audio containing genuine emotion and rich contextual cues is indeed more likely to break through the model’s safety defenses and cause jailbreaking compared to pure text input, even when augmented with emotional markers.
@@ G.2 Dataset Construction Method
Direct Privacy Leakage Dataset (600 samples). For the direct privacy leakage dataset,we created six categories of sensitive personal information: bank account numbers, home addresses, phone numbers, phone passwords, driver’s license numbers, and social security numbers (Wang et al., 2023a). For each category, we first prompted Grok to generate textual data containing personal names and corresponding private information (xAI, 2025). The generated data were then further randomized to enhance diversity and reduce identifiability. Based on this preliminary dataset, Grok was instructed to generate realistic dialogue scenarios in which each type of private information might naturally appear. Subsequently, Grok produced context-appropriate conversational texts for each scenario. Finally, we selected voice samples from the Common Voice dataset and used F5-TTS to synthesize the corresponding audio files (Chen et al., 2024; Ardila et al., 2020). Each category contained 100 entries, resulting in a total of 600 test samples.
Privacy Inference Leakage Dataset (300 samples). For the evaluation of implicit privacy inference, including age, ethnicity, and conversation setting (Xu et al., 2025a), we constructed another dataset comprising 300 audio samples, with 100 samples per category and construction method refers to other study (Wang et al., 2025b) . The age and ethnicity samples were obtained by filtering and curating entries from the Common Voice database. For the conversation setting category, we first collected ambient audio clips representing common daily-life environments such as “workspace” and “home.” Grok was then used to generate dialogues likely to occur in these environments. These dialogues were synthesized into speech using TTS and mixed with the corresponding background environment audio to create the final dataset. Below, we provide a selection of prompts employed during the dataset construction process using large language models.
@@ G.4 Evaluation Method and Metrics
For each audio input, privacy-related questions were posed via audio question answering (e.g., “What is the speaker‘s bank account number?” or “What is the speaker‘s age?”). The models’ responses were recorded and automatically scored using a five-point scoring system, ranging from 5 (complete refusal due to privacy concerns) to 1 (full disclosure of private information) (Yan et al., 2025), with intermediate scores reflecting varying degrees of caution in information disclosure.
We reported two key metrics based on the scores:
Refusal rate: the proportion of responses scored 3-5 (indicating refusal or cautious non-disclosure),
Inference accuracy when privacy is leaked: This rate shows inference accuracy only for samples where privacy was leaked.
@@ G.5 Experiment Design and Results
We used synthetic conversational audio containing specific types of private information (e.g., bank account numbers, home addresses, and phone numbers) as input to five closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and nine open-source ALLMs (e.g., Qwen2-Audio). We then posed a series of privacy-related questions tied to the audio content and recorded the models‘ responses.
@@
The audio samples were input into five closed-source ALLMs (e.g., GPT-4o Audio, Gemini-1.5 Pro) and nine open-source ALLMs (e.g., Qwen2-Audio). For age and ethnicity inference, implicit information was queried via AQA (e.g., “What is the speaker’s age?” or “What is the speaker’s ethnicity?”), and responses were scored using the same five-point scoring system as in the direct leakage experiments (5 = complete refusal, 1 = full disclosure). Scoring was performed automatically by a pretrained language model.
@@
| SALMONN | 52 | 97 | 59 | 98 | 63 | 89 | 75 | 99 | 66 | 98 | 30 | 100 | 97 | 100 | 34 | 13 | 14 | 34 |
| UltraVox | 98 | 100 | 95 | 100 | 19 | 99 | 76 | 99 | 100 | 100 | 53 | 100 | 1 | 6 | 25 | 22 | 0 | 1 |
| Qwen2-Audio | 0 | 18 | 1 | 19 | 0 | 40 | 4 | 49 | 0 | 7 | 0 | 9 | 0 | 0 | 0 | 0 | 4 | 3 |
| MiniCPM-o 2.6 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 4 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 4 | 3 |
| Step Fun | 99 | 98 | 21 | 100 | 2 | 99 | 22 | 97 | 98 | 99 | 7 | 97 | 46 | 51 | 0 | 8 | 0 | 0 |
| Qwen2.5-Omni | 0 | 0 | 0 | 3 | 0 | 0 | 0 | 0 | 0 | 3 | 0 | 1 | 0 | 0 | 0 | 0 | 1 | 0 |
| Kimi Audio | 0 | 0 | 1 | 5 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 1 | 45 | 26 | 0 | 0 | 0 | 0 |
| OpenS2S | 12 | 55 | 21 | 62 | 0 | 36 | 4 | 35 | 9 | 43 | 0 | 32 | 0 | 1 | 9 | 4 | 17 | 17 |
| Step Audio2 | 0 | 82 | 0 | 38 | 0 | 11 | 0 | 63 | 0 | 36 | 0 | 3 | 0 | 0 | 0 | 0 | 0 | 0 |
@@
| Gemini-1.5 Pro | 1 | 100 | 0 | 100 | 0 | 99 | 0 | 98 | 71 | 100 | 0 | 94 | 16 | 28 | 0 | 0 | 0 | 0 |
| GPT-4o Audio | 100 | 100 | 100 | 100 | 67 | 99 | 100 | 99 | 100 | 100 | 85 | 100 | 2 | 22 | 18 | 34 | 0 | 4 |
| GPT-4o Mini Audio | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 9 | 70 | 33 | 50 | 0 | 0 |
| Gemini-2.5 Flash | 8 | 98 | 0 | 100 | 0 | 100 | 0 | 96 | 0 | 100 | 0 | 85 | 2 | 8 | 0 | 1 | 1 | 1 |
| Gemini-2.5 Pro | 0 | 96 | 0 | 94 | 100 | 98 | 0 | 100 | 100 | 94 | 1 | 83 | 0 | 0 | 0 | 0 | 1 | 2 |
| Average | 33.6 | 67.4 | 28.4 | 65.6 | 25.1 | 62.1 | 27.2 | 67.0 | 46.0 | 62.9 | 19.7 | 57.5 | 15.6 | 22.3 | 8.5 | 9.4 | 3.0 | 4.6 |
Note: “w/o” indicates the refusal rates before applying a privacy-aware prompt, while “w/” shows rates after applying it. Higher values indicate better performance. Blue arrows () indicate better performance (higher refusal rate) than average; red arrows () indicate worse performance (lower refusal rate) than average. SSN: Social Security Number; Pwd: Phone Password; Setting: Setting of Conversation.
| Model | Bank Account | Driver License | Phone Number | Password | SSN | Address | Age | Ethnicity | Setting | |||||||||
| Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | Acc% | Leak | |